The music teacher at your school buys and sells fresh fruit to help raise funds to support the school concert band. One day the total order for bags of oranges and bags of apples cost $33.00 exactly. If 15 bags of fruit were purchased, how many bags contained oranges?
by the bag : oranges are $5.00 and apples are $2.00

Answer by rfer(16322)

5x+2(15-x)=33
5x+30-2x=33
3x=3
x=1------------
1 bag of oranges and
14 bags of apples
-----------------
(1*5)+(14*2)=33